OpenStudy (anonymous):

2.Indicate the equation of the given line in standard form. The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5) 3.The line that contains the point Q( 1, -2) and is parallel to the line whose equation is y - 4 = 2/3 (x - 3)

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

start by finding the slope of RS

ganeshie8 (ganeshie8):

R(-1, 6) and S(5, 5) slope = ?

ganeshie8 (ganeshie8):

can u find ? we're doing the first question btw..

OpenStudy (anonymous):

is it y = 6x - 6.5.

ganeshie8 (ganeshie8):

good try, but no.

ganeshie8 (ganeshie8):

slope = m = \(\large \dfrac{y_2 - y_1}{x_2 -x_1}\)

ganeshie8 (ganeshie8):

R(-1, 6) and S(5, 5) x1 y1 x2 y2

ganeshie8 (ganeshie8):

plug the values in slope formula, wat do u get ?

OpenStudy (anonymous):

uhhh

OpenStudy (anonymous):

im lost

ganeshie8 (ganeshie8):

slope = m = \(\large \dfrac{5-6}{5--1} = \dfrac{-1}{6}\)

OpenStudy (anonymous):

I see im writing it down and its getting confusing

ganeshie8 (ganeshie8):

just takes practice wid 2-3 problems to feel confident :)

OpenStudy (anonymous):

lol could i just get the answer that would save time since my school ends in 1 hour

OpenStudy (anonymous):

This is the last test

OpenStudy (anonymous):

if you don't beleeve we could screen share